OpenStudy (anonymous):

A parent died and left an estate to four children. One inherited 1/3, the second 1/3, the third 1/6. How much did the fourth inherit?

OpenStudy (anonymous):

Alright so in this problem your going to want to figure out what 1/3+1/3= ............

OpenStudy (anonymous):

Sorry I ment figure that out ans make it into 6ths then ad one 6th

OpenStudy (anonymous):

and*

OpenStudy (anonymous):

would the answer be 1/2

OpenStudy (anonymous):

Add 1/3+1/3=2/3 2/3=x/6

OpenStudy (anonymous):

3x/12

OpenStudy (anonymous):

Good, so what is your answer?

OpenStudy (anonymous):

This is how I thought of it: There are four children, so there are four fractions. Because one child is inheriting 1/6 of the estate all the other fractions must be expanded to suit this smaller fraction. 2/6 + 2/6 + 1/6 = 5/6. This is the known fraction of inheritance for the estate. 6/6 - 5/6 = 1/6. The forth child inherited one sixth of the estate

OpenStudy (anonymous):

I got it :) Thank you

OpenStudy (anonymous):

No problem!